# Adversarial Robustness in Two-Stage Learning-to-Defer: Algorithms and Guarantees

### v1

@@ 4 Adversarial Attacks on Two-Stage L2D
The two-stage L2D framework is designed to route queries to the most accurate agents, ensuring optimal decision-making Mao et al. (2023a, 2024d); Montreuil et al. (2024). Despite its effectiveness, we demonstrate that this framework is inherently vulnerable to adversarial attacks that exploit its reliance on the rejector function, a key component responsible for query allocation. Given this critical role, our analysis focuses on adversarial attacks and corresponding defenses targeting the rejector $r\in\mathcal{R}$, rather than individual agents. This focus is justified because adversarial defenses for specific agents can typically be deployed offline within the two-stage L2D setup. Moreover, evaluating the robustness of individual agents under adversarial conditions simplifies to selecting the most robust agent, whereas ensuring robustness at the system level constitutes a fundamentally different challenge.
Untargeted Attack:
@@
In classification, the goal of an untargeted attack is to find a perturbed input $x^{\prime}\in B_{p}(x,\gamma)$ that causes the classifier $h\in\mathcal{H}$ to misclassify the input (Goodfellow et al., 2014; Akhtar and Mian, 2018). Specifically, for a clean input $x\in\mathcal{X}$ where the classifier correctly predicts $h(x)=y$, the attacker aims to identify a perturbed input $x^{\prime}$ such that $h(x^{\prime})\neq y$.
@@
In the context of Learning-to-Defer, the attack extends beyond misclassification to compromising the decision allocation mechanism. Given an optimal agent $j^{\ast}\in\mathcal{A}$, the attacker aims to find an adversarial input $x^{\prime}\in B_{p}(x,\gamma)$ such that $r(x^{\prime})\neq j^{\ast}$, thereby forcing deferral to a suboptimal agent $j\in\mathcal{A}\setminus\{j^{\ast}\}$—incurring a higher loss. This leads to the following untargeted attack formulation:
Definition 2 (Untargeted Attack in L2D).
@@
Let $x^{\prime}\in B_{p}(x,\gamma)$ be an adversarial input, where $B_{p}(x,\gamma)$ denotes the $p$-norm ball of radius $\gamma$ centered at $x$. The untargeted attack that maximizes misallocation in L2D is formulated as follows:
$x^{\prime}=\operatorname*{arg\,sup}_{x^{\prime}\in B_{p}(x,\gamma)}\sum^{J}_{j=0}\tau_{j}(g(x),m(x),z)\Phi_{01}^{u}(r,x^{\prime},j).$
@@
Definition (2) characterizes an attack in which the adversary maximizes the aggregate discrepancy between the rejector’s allocation for the adversarial input $x^{\prime}$ and its original allocation, thereby forcing the system to defer to an unintended agent. The adversarial input $x^{\prime}$ is designed to maximize these surrogate losses. As a result, the adversarial attack increases the system’s overall loss, significantly degrading its performance. An illustration of this attack is provided in Appendix Figure 1.
Targeted Attack:
@@
Targeted attacks are often more impactful than untargeted ones, as they exploit specific system vulnerabilities to achieve precise adversarial goals (Akhtar and Mian, 2018; Chakraborty et al., 2021). For example, in autonomous driving classification, a targeted attack could deliberately misclassify a stop sign as a speed limit sign, potentially leading to hazardous consequences. A targeted attack exploits this asymmetry by forcing the classifier $h\in\mathcal{H}$ to predict a specific target class $y_{t}\in\mathcal{Y}$, potentially leading to harmful consequences.
@@
In the context of L2D, an attacker aims to manipulate the system into assigning a query $x^{\prime}\in B_{p}(x,\gamma)$ to a predetermined expert $j_{t}\in\mathcal{A}$ rather than the optimal expert $j^{\ast}\in\mathcal{A}$. This targeted attack objective can be formally expressed as follows:
Definition 3 (Targeted Attack in L2D).
@@
Let $x^{\prime}\in B_{p}(x,\gamma)$, where $B_{p}(x,\gamma)$ denotes the $p$-norm ball of radius $\gamma$ centered at $x$. The attack targeting the allocation of a query $x$ to the expert $j_{t}\in\mathcal{A}$ is defined as:
$x^{\prime}=\operatorname*{arg\,inf}_{x^{\prime}\in B_{p}(x,\gamma)}\tau_{j_{t}}(g(x),m(x),z)\Phi_{01}^{u}(r,x^{\prime},j_{t}).$
@@
The adversarial input $x^{\prime}$ minimizes the loss associated with the targeted agent $j_{t}$, thereby biasing the allocation process towards agent $j_{t}$ as described in Definition (3). For instance, an attacker may have an affiliated partner $j_{t}$ among the system’s agents. Suppose the system operates under a pay-per-query model—for example, a specialist doctor in a medical decision-making system or a third-party service provider in an AI-powered platform. By manipulating the allocation mechanism to systematically route more queries toward $j_{t}$, the attacker artificially inflates its workload, leading to unjustified financial gains. These gains may benefit both the attacker and the affiliated expert through direct financial compensation, revenue-sharing agreements, or other collusive incentives. An illustration of this attack is provided in Appendix Figure 2.
@@ 5 Adversarially Consistent Formulation for Two-Stage Learning-to-Defer
In this section, we introduce an adversarially consistent formulation of the two-stage L2D framework, ensuring robustness against attacks while preserving optimal query allocation.
@@ 5.1 Novel Two-stage Learning-to-Defer formulation
Adversarial True Deferral Loss:
@@
Lemma 4 (Adversarial True Deferral Loss).
@@
~$\widetilde{\ell}_{\text{def}}(r,g,m,z)$ $=\sum_{j=0}^{J}\tau_{j}(g(x),m(x),z)\widetilde{\ell}_{01}^{j}(r,x,j)+(1-J)\sum_{j=0}^{J}c_{j}(g(x),m_{j}(x),z).$
See Appendix D.1 for the proof of Lemma 4. The attacker’s objective is to compromise the allocation process by identifying perturbations $\delta_{j}$ that maximize the loss for each agent $j\in\mathcal{A}$. Importantly, the costs $c_{j}$ and $\tau_{j}$ are evaluated based on the clean input $x$, as the agents’ predictions remain unaffected by the perturbations (see Section 4).
Minimizing the adversarial true deferral loss in Lemma 4 is NP-hard (Zhang, 2002; Bartlett et al., 2006; Steinwart, 2007; Awasthi et al., 2022). Therefore, as in classification problems, we approximate this discontinuous loss using surrogates.
Adversarial Margin Deferral Surrogate Losses:
In the formulation of the adversarial true deferral loss (Lemma 4), discontinuities arise due to the indicator function in the loss definition. To approximate this discontinuity, we build on recent advancements in consistency theory for adversarially robust classification (Bao et al., 2021; Awasthi et al., 2022, 2023; Mao et al., 2023b) and propose a continuous, upper-bound surrogate family for the adversarial true deferral loss. Specifically, we define the $j$-th adversarial margin surrogate family $\widetilde{\Phi}^{\rho,u,j}_{01}(r,x,j)=\sup_{x_{j}^{\prime}\in B_{p}(x,\gamma)}\Psi^{u}\left(\sum_{j^{\prime}\neq j}\Psi_{\rho}\big{(}r(x_{j}^{\prime},j^{\prime})-r(x_{j}^{\prime},j)\big{)}\right)$ where $\Psi^{u}$ and $\Psi_{\rho}$ are defined in Equation (6). Building on this, we derive the adversarial margin deferral surrogate losses as:
Lemma 5 (Adversarial Margin Deferral Surrogate Losses).
@@
The proof is provided in Appendix D.2. One notable limitation of the adversarial margin deferral surrogate family is the non-convexity of the $j$-th adversarial margin surrogate loss family $\widetilde{\Phi}^{\rho,u,j}_{01}$, which poses significant challenges for efficient optimization.
Adversarial Smooth Deferral Surrogate Losses:
As detailed by Awasthi et al. (2023); Mao et al. (2023b), the non-convex adversarial margin surrogate family can be replaced with a smooth and convex approximation. To this end, we adapt their results and introduce the smooth adversarial surrogate family, denoted as $\widetilde{\Phi}^{\text{smth},u}_{01}:\mathcal{R}\times\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{+}$, which approximates the supremum term in Lemma 4. Crucially, $\widetilde{\Phi}^{\text{smth},u}_{01}$ acts as a convex, non-negative upper bound for the $j$-th adversarial margin surrogate family, such that $\widetilde{\Phi}_{01}^{\rho,u,j}\leq\widetilde{\Phi}_{01}^{\text{smth},u}$. We derive the smooth adversarial surrogate losses in Lemma 6.
Lemma 6 (Smooth Adversarial Surrogate Losses).
Let $x\in\mathcal{X}$ denote the clean input and hyperparameters $\rho>0$ and $\nu>0$. We define the smooth adversarial surrogate losses as:
@@
For completeness, the proof is provided in Appendix D.3. For $x\in\mathcal{X}$, define $\Delta_{r}(x,j,j^{\prime})=r(x,j)-r(x,j^{\prime})$, and let $\overline{\Delta}_{r}(x,j)$ denote the $J$-dimensional vector $\big{(}\Delta_{r}(x,j,0),\ldots,\Delta_{r}(x,j,j-1),\Delta_{r}(x,j,j+1),\ldots,\Delta_{r}(x,j,J)\big{)}$. The first term, $\Phi_{01}^{u}(r/\rho,x,j)$, corresponds to the multiclass surrogate losses modulated by the coefficient $\rho>0$. The second term incorporates adversarial evaluations $x_{j}^{\prime}\in B_{p}(x,\gamma)$ for each agent $j\in\mathcal{A}$, with a smooth adversarial component scaled by the coefficient $\nu>0$ (Awasthi et al., 2023; Mao et al., 2023b). The coefficients $(\rho,\nu)$ are typically selected through cross-validation to balance allocation performance and robustness against adversarial perturbations.
Using the smooth adversarial surrogate family from Lemma 6, we define the smooth adversarial deferral surrogate (SAD) family $\widetilde{\Phi}_{\text{def}}^{\text{smth},u}:\mathcal{R}\times\mathcal{G}\times\mathcal{M}\times\mathcal{Z}\to\mathbb{R}^{+}$, which is convex, non-negative, and serves as an upper bound for ~$\widetilde{\ell}_{\text{def}}$ by construction. The formal definition of our novel surrogates is given as:
Lemma 7 (SAD: Smooth Adversarial Deferral Surrogate Losses).
Let $x\in\mathcal{X}$ denote the clean input and $\tau_{j}$ the aggregated cost. Then, the smooth adversarial surrogate family (or SAD) $\widetilde{\Phi}^{\text{smth},u}_{\text{def}}$ is defined as:
$\widetilde{\Phi}^{\text{smth},u}_{\text{def}}(r,g,m,z)=\sum_{j=0}^{J}\tau_{j}(g(x),m(x),z)\widetilde{\Phi}_{01}^{\text{smth},u}(r,x,j).$
The proof is provided in Appendix D.4. While the smooth adversarial deferral surrogate family (SAD) provides a smooth and computationally efficient approximation of the adversarial true deferral loss, the question of its consistency remains a critical consideration (Zhang, 2002; Bartlett et al., 2006).
@@ 5.2 Theoretical Guarantees
To establish the theoretical foundations of SAD, we prove the Bayes-consistency and $(\mathcal{R},\mathcal{G})$-consistency of the adversarial margin deferral surrogate family $\widetilde{\Phi}^{\rho,u}_{\text{def}}$ (Lemma 5). Furthermore, we demonstrate that these guarantees naturally extend to a regularized empirical formulation of SAD, referred to as SARD.
$(\mathcal{R},\mathcal{G})$-consistency bounds of $\widetilde{\Phi}^{\rho,u}_{\text{def}}$:
A key foundational step involves demonstrating the $\mathcal{R}$-consistency of the $j$-th adversarial margin surrogate losses $\widetilde{\Phi}^{\rho,u,j}_{01}$, under the assumption that $\mathcal{R}$ is symmetric and that there exists a rejector $r\in\mathcal{R}$ that is locally $\rho$-consistent.
Definition 8 (Locally $\rho$-consistent).
@@
where $\rho>0$, $i\neq j\in\mathcal{A}$, and $x^{\prime}\in B_{p}(x,\gamma)$. Additionally, for any $x^{\prime}\in B_{p}(x,\gamma)$, the set $\{r(x^{\prime},j):j\in\mathcal{A}\}$ preserves the same ordering as for $x$.
As shown in Awasthi et al. (2022); Mao et al. (2023b); Awasthi et al. (2023), commonly used hypothesis sets, such as linear models, neural networks, and the set of all measurable functions, are locally $\rho$-consistent for some $\rho>0$. Consequently, the guarantees established in Lemma 9 are general and broadly applicable across diverse practical settings. The proof of Lemma 9 is deferred to Appendix D.5.
Lemma 9 ($\mathcal{R}$-consistency bounds for $\widetilde{\Phi}^{\rho,u,j}_{01}$).
Assume $\mathcal{R}$ is symmetric and locally $\rho$-consistent. Then, for the agent set $\mathcal{A}$, any hypothesis $r\in\mathcal{R}$, and any distribution $\mathcal{P}$ with probabilities $p=(p_{0},\cdots,p_{J})\in\Delta^{|\mathcal{A}|}$, the following inequality holds:
@@
Lemma 9 establishes the consistency of the $j$-th adversarial margin surrogate family $\widetilde{\Phi}^{\rho,u,j}_{01}$ for probabilities $p_{j}\in\Delta^{|\mathcal{A}|}$, explicitly incorporating adversarial inputs defined for each $j\in\mathcal{A}$. This result distinguishes our contribution from prior works (Mao et al., 2023b; Awasthi et al., 2023, 2022), which do not address adversarial inputs at the level of the distribution (dependent on $j$). By addressing this limitation, Lemma 9 provides a critical theoretical guarantee, demonstrating that the $j$-th adversarial margin surrogate family $\widetilde{\Phi}^{\rho,u,j}_{01}$ aligns with the adversarial loss ~$\widetilde{\ell}_{01}^{j}$ under the specified assumptions.
Building on the foundational result of Lemma 9, we prove the Bayes and $(\mathcal{R},\mathcal{G})$-consistency of the adversarial margin deferral surrogate losses. The proof of Theorem 10 is provided in Appendix D.6.
Theorem 10 ($(\mathcal{R},\mathcal{G})$-consistency bounds of $\widetilde{\Phi}^{\rho,u}_{\text{def}}$).

### v2

@@ 4 Novel Adversarial Attacks on L2D
The two-stage L2D framework is designed to route queries to the most accurate agents, ensuring optimal decision-making (Narasimhan et al., 2022; Mao et al., 2023a, 2024c; Montreuil et al., 2024). Despite its effectiveness, we show that this framework is inherently vulnerable to adversarial attacks that exploit its reliance on the rejector function—a key component responsible for query allocation. Given this critical role, our analysis concentrates on adversarial attacks and corresponding defenses targeting the rejector $r\in\mathcal{R}$, rather than individual agents. This emphasis is warranted by the fact that adversarial defenses for specific agents can typically be deployed offline within the two-stage L2D setup, whereas the rejector operates as a router at inference time, serving as the first layer in the architecture to direct queries to the appropriate agent. Moreover, evaluating the robustness of individual agents under adversarial conditions reduces to selecting the most robust agent, whereas ensuring robustness at the system level constitutes a fundamentally different challenge.
Untargeted Attack.
@@
In classification, the goal of an untargeted attack is to find a perturbed input $x^{\prime}\in B_{p}(x,\gamma)$ that causes the classifier $h\in\mathcal{H}$ to misclassify the input (Goodfellow et al., 2014; Akhtar & Mian, 2018). Specifically, for a clean input $x\in\mathcal{X}$ such that the classifier correctly predicts $h(x)=y$, the attacker seeks a perturbed input $x^{\prime}$ such that $h(x^{\prime})\neq y$.
@@
In the context of Learning-to-Defer, the attack extends beyond misclassification to compromising the decision allocation mechanism. Given an optimal agent $j^{\ast}\in\mathcal{A}$, the attacker aims to find an adversarial input $x^{\prime}\in B_{p}(x,\gamma)$ such that $r(x^{\prime})\neq j^{\ast}$, thereby resulting in a deferral to a suboptimal agent $j\in\mathcal{A}\setminus\{j^{\ast}\}$, leading to an increased loss. This gives rise to the following untargeted attack formulation:
Definition 4.1 (Untargeted Attack in L2D).
@@
Let $x^{\prime}\in B_{p}(x,\gamma)$ be an adversarial input, where $B_{p}(x,\gamma)$ denotes the $p$-norm ball of radius $\gamma$ centered at $x$. The untargeted attack that maximizes misallocation in L2D is defined as:
$x^{\prime}=\underset{x^{\prime}\in B_{p}(x,\gamma)}{\arg\max}\sum_{j=0}^{J}\tau_{j}(g(x),m(x),z)\Phi_{01}^{u}(r,x^{\prime},j)$
@@
Definition 4.1 characterizes an attack in which the adversary maximizes the overall surrogate loss incurred by misallocation, thereby causing the system to defer to an unintended agent. The adversarial input $x^{\prime}$ is designed to maximize these surrogate losses. As a result, the attack increases the system’s overall loss, leading to a significant degradation in performance. An illustration of this attack is provided in Appendix Figure 2.
Targeted Attack.
@@
Targeted attacks are often more impactful than untargeted ones, as they exploit specific system vulnerabilities to achieve precise adversarial goals (Akhtar & Mian, 2018; Chakraborty et al., 2021). For example, in autonomous driving, a targeted attack could deliberately misclassify a stop sign as a speed limit sign, potentially leading to hazardous consequences. Such attacks exploit asymmetries in the task by forcing the classifier $h\in\mathcal{H}$ to predict a specific target class $y_{t}\in\mathcal{Y}$, thereby amplifying the risk of incorrect or dangerous decisions.
@@
In the context of L2D, the attacker seeks to manipulate the system to assign a query $x^{\prime}\in B_{p}(x,\gamma)$ to a predetermined expert $j_{t}\in\mathcal{A}$, rather than the optimal expert $j^{\ast}\in\mathcal{A}$. This targeted attack can be formally expressed as follows:
Definition 4.2 (Targeted Attack in L2D).
@@
Let $x^{\prime}\in B_{p}(x,\gamma)$, where $B_{p}(x,\gamma)$ denotes the $p$-norm ball of radius $\gamma$ centered at $x$. The targeted attack that biases the allocation of a query $x$ toward a predetermined expert $j_{t}\in\mathcal{A}$ is defined as
$x^{\prime}=\underset{x^{\prime}\in B_{p}(x,\gamma)}{\arg\min}\,\tau_{j_{t}}(g(x),m(x),z)\Phi_{01}^{u}(r,x^{\prime},j_{t}).$
The adversarial input $x^{\prime}$ minimizes the surrogate loss associated with the target agent $j_{t}$, thereby biasing the allocation process toward agent $j_{t}$ (Definition 4.2).
@@
For instance, an attacker may have an affiliated partner $j_{t}$ among the system’s agents. Suppose the system operates under a pay-per-query model—such as a specialist doctor in a medical decision-making system or a third-party service provider in an AI platform. By manipulating the allocation mechanism to systematically route more queries to $j_{t}$, the attacker artificially inflates the agent’s workload, resulting in unjustified financial gains shared between the attacker and the affiliated expert through direct payments, revenue-sharing, or other collusive arrangements. An illustration of this attack is provided in Appendix Figure 3.
@@ 5 Adversarially Consistent Formulation for Two-Stage Learning-to-Defer
In this section, we introduce an adversarially consistent formulation of the two-stage L2D framework that ensures robustness against attacks while preserving optimal query allocation.
@@ 5.1 Novel Two-stage Learning-to-Defer Formulation
Adversarial True Deferral Loss.
@@
Definition 5.1 ($j$-th Adversarial True Multiclass Loss).
We define the $j$-th adversarial true multiclass loss as
~$\widetilde{\ell}_{01}^{j}(r,x,j)=\sup_{x_{j}^{\prime}\in B_{p}(x,\gamma)}\ell_{01}(r(x_{j}^{\prime}),j),$
which captures the worst-case misclassification loss when deferring to agent $j$ under adversarial conditions. The adversarial true deferral loss is defined in Lemma 5.2.
Lemma 5.2 (Adversarial True Deferral Loss).
@@
~$\widetilde{\ell}_{\text{def}}(r,g,m,z)$ $=\sum_{j=0}^{J}\tau_{j}(g(x),m(x),z)\widetilde{\ell}_{01}^{j}(r,x,j)$
$+(1-J)\sum_{j=0}^{J}c_{j}(g(x),m_{j}(x),z).$
See Appendix D.1 for the proof of Lemma 5.2. The attacker’s objective is to compromise the allocation process by identifying perturbations $\delta_{j}$ that maximize the loss for each agent $j\in\mathcal{A}$. Notably, the costs $c_{j}$ and $\tau_{j}$ are evaluated on the clean input $x$, as the agents’ predictions remain unaffected by the perturbations (see Section 4).
Remark 5.3.
Simply applying adversarial evaluation to the entire two-stage loss (Definition 3.3) does not account for the worst-case scenario captured in Lemma 5.2. This point is further discussed in Appendix D.1.
As with standard classification losses, minimizing the adversarial true deferral loss in Lemma 5.2 is NP-hard (Zhang, 2002; Bartlett et al., 2006; Steinwart, 2007; Awasthi et al., 2022). Therefore, as in classification problems, we approximate this discontinuous loss using surrogates.
Adversarial Margin Deferral Surrogate Losses.
In the formulation of the adversarial true deferral loss (Lemma 5.2), discontinuities arise from the use of the indicator function in its definition. To approximate this discontinuity, we build on recent advancements in consistency theory for adversarially robust classification (Bao et al., 2020; Awasthi et al., 2022, 2023; Mao et al., 2023b) and propose a continuous, upper-bound surrogate family to approximate it.
Definition 5.4 ($j$-th Adversarial Margin Surrogates).
We define the $j$-th adversarial margin surrogate family as
[…]
where $\Psi^{u}$ and $\Psi_{\rho}$ are defined in Equation (3.2). Based on this definition, we derive the adversarial margin deferral surrogate losses as:
Lemma 5.5 (Adversarial Margin Deferral Surrogate Losses).
@@
The proof is provided in Appendix D.2. A key limitation of the adversarial margin deferral surrogate family lies in the non-convexity of the $j$-th adversarial margin surrogate loss $\widetilde{\Phi}^{\rho,u,j}_{01}$, induced by the piecewise nature of $\Psi_{\rho}(v)$, which complicates efficient optimization.
Adversarial Smooth Deferral Surrogate Losses.
Prior work by Awasthi et al. (2023); Mao et al. (2023b) shows that the non-convex adversarial margin surrogate family can be approximated by a smooth and convex surrogate. Building on their results, we introduce the smooth adversarial surrogate family, denoted as $\widetilde{\Phi}^{\text{smth},u}_{01}:\mathcal{R}\times\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{+}$, which approximates the supremum term in Lemma 5.2. Importantly, $\widetilde{\Phi}^{\text{smth},u}_{01}$ serves as a convex, non-negative upper bound for the $j$-th adversarial margin surrogate family, satisfying $\widetilde{\Phi}_{01}^{\rho,u,j}\leq\widetilde{\Phi}_{01}^{\text{smth},u}$. We derive this surrogate in Lemma 5.6.
Lemma 5.6 (Smooth Adversarial Surrogate Losses).
Let $x\in\mathcal{X}$ denote the clean input and hyperparameters $\rho>0$ and $\nu>0$. The smooth adversarial surrogate losses are defined as:
$\widetilde{\Phi}_{01}^{\text{smth},u}(r,x,j)=\Phi_{01}^{u}$ $(\frac{r}{\rho},x,j)$
@@
The proof is provided in Appendix D.3. For $x\in\mathcal{X}$, define $\Delta_{r}(x,j,j^{\prime})=r(x,j)-r(x,j^{\prime})$, and let $\overline{\Delta}_{r}(x,j)\in\mathbb{R}^{J}$ denote the vector of pairwise differences $\big{(}\Delta_{r}(x,j,0),\ldots,\Delta_{r}(x,j,j-1),\Delta_{r}(x,j,j+1),\ldots,\Delta_{r}(x,j,J)\big{)}$. The first term $\Phi_{01}^{u}\left(\frac{r}{\rho},x,j\right)$ corresponds to the multiclass surrogate losses modulated by the coefficient $\rho$. The second term introduces a smooth adversarial component over perturbations $x_{j}^{\prime}\in B_{p}(x,\gamma)$, scaled by the coefficient $\nu$ (Awasthi et al., 2023; Mao et al., 2023b). The hyperparameters $(\rho,\nu)$ are typically selected via cross-validation to balance allocation performance and robustness.
Using this surrogate, we define the smooth adversarial deferral surrogate (SAD) family $\widetilde{\Phi}_{\text{def}}^{\text{smth},u}:\mathcal{R}\times\mathcal{G}\times\mathcal{M}\times\mathcal{Z}\to\mathbb{R}^{+}$, which is convex, non-negative, and upper-bounds ~$\widetilde{\ell}_{\text{def}}$ by construction. Its formal definition is as follows:
Lemma 5.7 (SAD: Smooth Adversarial Deferral Surrogate Losses).
Let $x\in\mathcal{X}$ denote the clean input and $\tau_{j}$ the aggregated cost. Then, the smooth adversarial surrogate family (SAD) $\widetilde{\Phi}^{\text{smth},u}_{\text{def}}$ is defined as:
$\widetilde{\Phi}^{\text{smth},u}_{\text{def}}(r,g,m,z)=\sum_{j=0}^{J}\tau_{j}(g(x),m(x),z)\widetilde{\Phi}_{01}^{\text{smth},u}(r,x,j)$
The proof is provided in Appendix D.4. While the SAD family offers a smooth and computationally efficient approximation of the adversarial true deferral loss, the question of its consistency remains a central concern.
@@ 5.2 Theoretical Guarantees
To establish the theoretical foundations of SAD, we first prove that the family of adversarial margin deferral surrogates $\widetilde{\Phi}^{\rho,u}_{\text{def}}$ is Bayes-consistent and $(\mathcal{R},\mathcal{G})$-consistent (Lemma 5.5). Wang & Scott (2023) introduced sufficient conditions for a surrogate loss to satisfy Bayes-consistency. These conditions—Gamma-PD (Definition 3.1) and Phi-NDZ (Definition 3.2)—are jointly sufficient to apply their general Theorem 2.6.
While our surrogate satisfies the Gamma-PD property, it fails to meet the Phi-NDZ criterion because it is not differentiable on all of $\mathbb{R}$. As a consequence, Theorem 2.6 from Wang & Scott (2023) cannot be invoked directly to establish classification-calibration. This limitation motivates a dedicated analysis of the consistency properties of our surrogates. We show that both Bayes- and $(\mathcal{R},\mathcal{G})$-consistency can nonetheless be established for the class $\widetilde{\Phi}^{\rho,u}_{\text{def}}$, and that these guarantees extend to the regularized empirical variant used in our algorithmic implementation, referred to as SARD.
$(\mathcal{R},\mathcal{G})$-consistency bounds of $\widetilde{\Phi}^{\rho,u}_{\text{def}}$.
A key step toward establishing theoretical guarantees is to demonstrate the $\mathcal{R}$-consistency of the $j$-th adversarial margin surrogate losses $\widetilde{\Phi}^{\rho,u,j}_{01}$, under the assumption that $\mathcal{R}$ is symmetric and that there exists a rejector $r\in\mathcal{R}$ that is locally $\rho$-consistent.
Definition 5.8 (Locally $\rho$-consistent).
@@
where $\rho>0$, $i\neq j\in\mathcal{A}$, and $x^{\prime}\in B_{p}(x,\gamma)$. Moreover, the ordering of the values $\{r(x^{\prime},j)\}$ is preserved with respect to $\{r(x,j)\}$ for all $x^{\prime}\in B_{p}(x,\gamma)$.
As shown by Awasthi et al. (2022); Mao et al. (2023b); Awasthi et al. (2023), common hypothesis classes—including linear models, neural networks, and the set of all measurable functions—are locally $\rho$-consistent for some $\rho>0$. Consequently, the guarantees established in Lemma 5.9 are broadly applicable in practice. See Appendix D.5 for the proof.
Lemma 5.9 ($\mathcal{R}$-consistency bounds for $\widetilde{\Phi}^{\rho,u,j}_{01}$).
Assume $\mathcal{R}$ is symmetric and locally $\rho$-consistent. Then, for the agent set $\mathcal{A}$, any hypothesis $r\in\mathcal{R}$, and any distribution $\mathcal{P}$ with probabilities $p=(p_{0},\ldots,p_{J})\in\Delta^{|\mathcal{A}|}$, the following inequality holds:
@@
Lemma 5.9 establishes the consistency of the $j$-th adversarial margin surrogate family $\widetilde{\Phi}^{\rho,u,j}_{01}$ under adversarial perturbations defined for each $j\in\mathcal{A}$, weighted by distributional probabilities $p_{j}\in\Delta^{|\mathcal{A}|}$. This result sets our work apart from prior approaches (Mao et al., 2023b; Awasthi et al., 2023, 2022), which do not account for adversarial inputs at the per-agent distributional level. By addressing this limitation, Lemma 5.9 provides a crucial theoretical guarantee: it shows that the surrogate losses $\widetilde{\Phi}^{\rho,u,j}_{01}$ align with the true adversarial losses ~$\widetilde{\ell}_{01}^{j}$ under appropriate conditions.
Building on this result, we establish the Bayes- and $(\mathcal{R},\mathcal{G})$-consistency of the adversarial margin deferral surrogate losses. The proof of Theorem 5.10 is provided in Appendix D.6.
Theorem 5.10 ($(\mathcal{R},\mathcal{G})$-consistency bounds of $\widetilde{\Phi}^{\rho,u}_{\text{def}}$).